# arXiv:2603.01896v1[cs.SE] 02 Mar 2026

### v1

@@ Abstract
Can LLM agents explore codebases and reason about code semantics without executing the code? We study this capability, which we call agentic code reasoning, and introduce semi-formal reasoning: a structured prompting methodology that requires agents to construct explicit premises, trace execution paths, and derive formal conclusions. Unlike unstructured chain-of-thought, semi-formal reasoning acts as a certificate: the agent cannot skip cases or make unsupported claims. We evaluate across three tasks (patch equivalence verification, fault localization, and code question answering) and show that semi-formal reasoning consistently improves accuracy on all of them. For patch equivalence, accuracy improves from 78% to 88% on curated examples and reaches 93% on real-world agent-generated patches, approaching the reliability needed for execution-free RL reward signals. For code question answering on RubberDuckBench rubberduckbench, semi-formal reasoning achieves 87% accuracy. For fault localization on Defects4J just2014defects4j, semi-formal reasoning improves Top-5 accuracy by 5 percentage points over standard reasoning. These results demonstrate that structured agentic reasoning enables meaningful semantic code analysis without execution, opening practical applications in RL training pipelines, code review, and static program analysis.
@@ 4.2 Fault Localization
For matching, we compare overlaps against buggy file regions — the lines in the patch, grouped per hunk into (file, min_deleted, max_deleted) ranges for deletions, and (file, insert_pos, insert_pos) for insertions. A prediction matches a region if their ranges overlap: pred_start <= region_end and pred_end >= region_start. This works for all patch types including deletion-only cases, which can be an issue with matching based on line number in the ground truth patch. With this matching, a bug is "solved at Top-N" only when all ground-truth hunks/lines/regions are covered by predictions in positions 1..N. (This deviates slightly from prior work.)
@@
Table 4: Fault localization accuracy on Defects4J (Opus-4.5)
Mode Exploration Top-1 Top-3 Top-5
Standard Single-shot 36.1% 55.6% 55.6%
Semi-formal Single-shot 41.7% 58.3% 63.9%
Standard Agentic 46.5% 60.5% 60.5%
Semi-formal Agentic 53.5% 67.4% 72.1%
Semi-formal reasoning improves Top-5 accuracy by 8–12pp across both exploration modes: +8pp for single-shot and +12pp for agentic. Agentic exploration with semi-formal reasoning achieves the best overall accuracy at 72.1% Top-5, outperforming single-shot semi-formal (63.9%).
@@
Table 5 shows semi-formal reasoning improves accuracy by +5pp Top-5 over standard, confirming that semi-formal reasoning consistently outperforms standard mode across dataset sizes.
Table 5: Agentic fault localization on 100-bug sample, both Sonnet 4.5 and Opus 4.5
Model Mode Top-1 Top-3 Top-5
Sonnet-4.5 Standard 20.0% 30.0% 31.1%
Sonnet-4.5 Semi-formal 16.7% 25.6% 30.0%
Opus-4.5 Standard 30.0% 42.2% 43.3%
Opus-4.5 Semi-formal 34.4% 47.8% 47.8%
@@ 6 Conclusion and Future Work
  • For patch equivalence, we achieve 93% verification accuracy on real-world patches, enabling execution-free feedback for RL training pipelines.
  • For code question answering, semi-formal reasoning achieves 87% accuracy on RubberDuckBench, demonstrating effective semantic understanding of complex codebases.

### v2

# Shubham Ugare at

@@ Abstract
Can LLM agents explore codebases and reason about code semantics without executing the code? We study this capability, which we call agentic code reasoning, and introduce semi-formal reasoning: a structured prompting methodology that requires agents to construct explicit premises, trace execution paths, and derive formal conclusions. Unlike unstructured chain-of-thought, semi-formal reasoning acts as a certificate: the agent cannot skip cases or make unsupported claims. We evaluate across three tasks (patch equivalence verification, fault localization, and code question answering) and show that semi-formal reasoning consistently improves accuracy on all of them. For patch equivalence, accuracy improves from 78% to 88% on curated examples and reaches 93% on real-world agent-generated patches, approaching the reliability needed for execution-free RL reward signals. For code question answering on RubberDuckBench rubberduckbench, semi-formal reasoning achieves 87% accuracy, a 9 percentage point gain over standard agentic reasoning. For fault localization on Defects4J just2014defects4j, semi-formal reasoning improves Top-5 accuracy by 5 percentage points over standard reasoning. These results demonstrate that structured agentic reasoning enables meaningful semantic code analysis without execution, opening practical applications in RL training pipelines, code review, and static program analysis.
@@ 4.2 Fault Localization
For matching, we compare overlaps against buggy file regions: the lines in the patch, grouped per hunk into (file, min_deleted, max_deleted) ranges for deletions, and (file, insert_pos, insert_pos) for insertions. A prediction matches a region if their ranges overlap: pred_start $\leq$ region_end and pred_end $\geq$ region_start. This works for all patch types including deletion-only cases, which can be an issue with matching based on line number in the ground truth patch. We report two variants: All, where a bug is solved at Top-N only when every ground-truth hunk is covered by predictions in positions 1..N, and Any, where at least one ground-truth hunk must be covered. The All metric is stricter and penalizes multi-hunk bugs; the Any metric better reflects partial localization success.
@@
Table 4: Fault localization accuracy on Defects4J (Opus-4.5). “All” requires every ground-truth hunk to be covered; “Any” requires at least one.
All Any
Mode Exploration Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
Standard Single-shot 36.1% 55.6% 55.6% 47.2% 69.4% 69.4%
Semi-formal Single-shot 41.7% 58.3% 63.9% 55.6% 75.0% 77.8%
Standard Agentic 46.5% 60.5% 60.5% 65.1% 79.1% 81.4%
Semi-formal Agentic 53.5% 67.4% 72.1% 72.1% 83.7% 88.4%
Semi-formal reasoning improves accuracy across both metrics and exploration modes. Under the stricter All metric, Top-5 accuracy improves by +8pp for single-shot and +12pp for agentic. Under the Any metric, the gains are similar: +8pp for single-shot and +7pp for agentic. Agentic exploration with semi-formal reasoning achieves the best overall accuracy at 72.1% Top-5 (All) and 88.4% Top-5 (Any).
@@
Table 5 shows semi-formal reasoning improves accuracy by +5pp Top-5 (All) over standard, confirming that semi-formal reasoning consistently outperforms standard mode across dataset sizes. The Any metric, which requires at least one ground-truth hunk to be covered, shows the same trend with higher absolute numbers.
Table 5: Agentic fault localization on 100-bug sample. “All” requires every ground-truth hunk to be covered; “Any” requires at least one.
All Any
Model Mode Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
Sonnet-4.5 Standard 20.0% 30.0% 31.1% 35.6% 51.1% 51.1%
Sonnet-4.5 Semi-formal 16.7% 25.6% 30.0% 30.0% 43.3% 50.0%
Opus-4.5 Standard 30.0% 42.2% 43.3% 53.3% 64.4% 65.6%
Opus-4.5 Semi-formal 34.4% 47.8% 47.8% 58.9% 68.9% 68.9%
@@ 6 Conclusion and Future Work
  • For patch equivalence, we achieve 93% verification accuracy on real-world patches, a 7 percentage point improvement over single-shot baselines, enabling execution-free feedback for RL training pipelines.
  • For code question answering, semi-formal reasoning achieves 87% accuracy on RubberDuckBench, a 9 percentage point gain over standard agentic reasoning.